### v1

# Relative energy constraint in Boltzmann’s surface entropy eliminates thermodynamic inconsistencies and allows negative absolute temperatures

Abstract
Over the past century, an intense debate in statistical mechanics has been about the correctness of Boltzmann’s surface entropy versus Gibbs’ volume entropy, for isolated systems. Both entropies make significantly different predictions for systems with few degrees of freedom. Even in the thermodynamic limit, they can disagree—while Boltzmann entropy allows negative absolute temperatures to exist, Gibbs entropy precludes such a possibility. Here, we show that modifying Boltzmann’s entropy via a relative energy tolerance eliminates thermodynamic inconsistencies in several model systems with unbounded energy spectra by ensuring positive, finite temperatures. Concomitantly, the proposed entropy allows for negative temperatures in systems with bounded spectra and closely matches canonical ensemble predictions. This work conclusively remedies the prevalent deficiencies of the Gibbs and Boltzmann entropy formulations and paves the way for the use of the modified Boltzmann entropy in the microcanonical ensemble, allowing negative temperatures to exist.
The microcanonical ensemble in statistical mechanics, originally proposed by Gibbs over a century ago [1], continues to receive a significant extent of attention [2, 3, 4, 5, 6]. In the physics literature, there exist two competing entropy definitions in the microcanonical ensemble: Gibbs’ volume entropy (also referred to as the Hertz entropy) [1, 7, 8, 9, 10], based on the phase space volume and Boltzmann’s surface entropy [11, 12] (also referred to as the Boltzmann-Planck entropy), based on the phase space density. Even though Boltzmann’s surface entropy dominated textbook discourses [13, 14, 15, 16, 17], many studies favored the use of the Gibbs entropy [18, 8, 19, 9, 10, 5]. About a decade ago, Dunkel and Hilbert indicated that Boltzmann’s surface entropy leads to violations of equipartition and unphysical results, i.e., negative/infinite temperatures, for some model systems with unbounded spectra [5]. Subsequently, several articles appeared that pointed out that Gibbs entropy suffers from serious objections such as the consideration of all states with energy lower than energy $E$, rather than only the states at the desired energy, and the inability to allow for negative temperatures in systems with non-monotonically increasing density of states [20, 12, 21, 22, 23, 24, 6].
Although the Boltzmann and Gibbs entropies usually agree for very large systems, they differ significantly and lead to widely varying temperatures for isolated systems with few degrees of freedom [25]. (In some cases, even in the thermodynamic limit, e.g., the case of a two-state system (see below), they can disagree.) Note that the principles of statistical mechanics can be applied to systems with few degrees of freedom, as the statistical averaging is carried out over the classical phase space or the available quantum states, rather than over the degrees of freedom [26, 27, 28, 6]. Thus, even quantities like the temperature of an isolated quantum state [29, 30, 31] or the entropy of a few-body system are well defined [32, 33].
In light of this fact, that both entropy definitions have shortcomings for small systems [22] is particularly unnerving. Indeed, the lack of a universal microcanonical entropy formulation not only adversely impacts the fundamental physical understanding of the thermodynamics of small systems [34], but also could impede technological progress in the future, wherein systems of few atoms, molecules, or electrons are increasingly becoming physically realizable [35]. In this work, we show that the consideration of a relative energy tolerance in the microcanonical ensemble eliminates all thermodynamic inconsistencies put forth in previous work pertaining to systems with unbounded spectra, while at the same time allowing negative temperatures in systems with bounded spectra.
Formulation of the relative-energy constrained Boltzmann entropy. Considering the microcanonical ($NVE$) ensemble with fixed number of particles $N$, volume $V$, and energy $E$, the Gibbs entropy ($S_{G}$) is defined as $S_{G}=k_{B}\ln\Omega$, where $\Omega$ denotes the volume of the classical phase space or the number of quantum states in the region $E^{\prime}<E$ (Figure 1a). For a classical system, this can be written as
[…]
where $\Theta$ is the Heaviside step function, $\Theta(x)=\begin{cases}1;\;\;\;x>0\\ 0;\;\;\;x\leq 0\end{cases}$, $h$ is Planck’s constant, $d$ is the dimensionality of the system, $\mathrm{\textbf{r}}^{N}$ denotes the positions of the particles, and $\mathrm{\textbf{p}}^{N}$ denotes their momenta. Alternatively, for a quantum-mechanical system, we can write
$\Omega=\mathrm{Tr}[\Theta(E-H)]$ (1)
where $\mathrm{Tr}$ denotes the trace, i.e., the sum over quantum states in a Hilbert space. In contrast, the Boltzmann entropy ($S_{B}$) is conventionally defined as $S_{B}=k_{B}\ln\omega$, where
[…]
or
$\omega=\mathrm{Tr}\left[\delta\left(\frac{H-E}{\epsilon}\right)\right]$ (2)
and $\epsilon$ is an absolute energy tolerance, as visualized in Figure 1a. Physically, the tolerance appears because no system is perfectly isolated, and some amount of perturbation or broadening in the energy occurs due to the system’s interactions with its surroundings (Figure 1b).
From Eq. (1) and (2), one can infer that
$\omega=\epsilon\frac{\partial\Omega}{\partial E}$ (3)
Even for a discrete spectrum, although direct enumeration is often used as a simplification [36], Eq. (3) is preferred [14, 37], as it allows the correct definition of entropy for nondegenerate quantum systems [38, 39]. Typically, $\epsilon$ has been interpreted to be a fixed constant in Eq. (3), and this has led to inconsistent results for simple systems with unbounded energy levels, e.g., negative temperatures for an ideal gas with one degree of freedom or infinite temperature for a quantum harmonic oscillator and an ideal gas with two degrees of freedom. We suggest here that instead of using an absolute energy tolerance, the energy tolerance in the microcanonical ensemble should be defined relative to the system’s energy (Figure 1c). Indeed, a system with 1 J of energy and with 1 eV = 1.602 $\times$ 10-19 J of energy should not use the same tolerance in the microcanonical ensemble. This is because the higher the energy of a system, the greater its interaction with its surroundings, causing more broadening or perturbation in its energy. Thus, we propose that the constraint inside the delta function should be expressed in terms of the relative error of $H_{r}=\frac{H}{E}$ from 1. In terms of the phase space integrals or Hilbert space summations, doing so amounts to applying a constraint on $H$ without using up a degree of freedom. Thus, classically, one should write
[…]
whereas for a quantum-mechanical system, we should write
$\omega_{r}=\text{Tr}\left[\delta\left(\frac{H_{r}-1}{\epsilon_{r}}\right)\right]$ (5)
where $\epsilon_{r}$ is a relative error tolerance. In Eqs. (4) and (5), we have used the subscript $r$ to denote the use of the relative deviation from the desired energy. It can be shown that $\omega_{r}=\epsilon_{r}E\frac{\partial\Omega}{\partial E}$, as indicated in Figure 1c. Note that there is some precedent in the literature for adopting an energy-dependent tolerance [39, 38].
Figure 1: Motivation and explanation of the relative-energy-constrained Boltzmann entropy. (a) For a density of states $n(E)$, the Gibbs entropy is defined by the logarithm of the area under the curve, i.e., $\Omega$, while the Boltzmann entropy is defined by the logarithm of the change in area due to a fixed energy perturbation $\epsilon$, i.e., $\frac{\partial\Omega}{\partial E}\epsilon$. (b) Interaction of a system with its surroundings leads to smearing of the system’s energy $E$ by a perturbation $\epsilon$. (c) The proposed entropy considers an energy-dependent energy perturbation, given as $\epsilon_{r}E$.
We now apply the new approach for calculating the microcanonical relative-energy-constrained Boltzmann entropy, $S_{B_{r}}=k_{B}\ln\omega_{r}$, and the corresponding temperature
$\frac{1}{T_{B_{r}}}=\frac{\partial S_{B_{r}}}{\partial E}$ (6)
to several example systems. One may think that converting an absolute energy constraint to a relative energy constraint would not affect the calculation of entropy significantly. However, as seen in Table 1, we conclusively show that using a relative constraint preserves the degrees of freedom of a system and thus leads to thermodynamically consistent behavior for various isolated systems with few degrees of freedom.
Table 1: Comparison of the predictions of the Gibbs, conventional Boltzmann, and proposed modified Boltzmann entropies in the microcanonical ensemble with the corresponding predictions in the canonical ensemble. Only the relative-energy-constrained Boltzmann entropy leads to predictions consistent with the canonical ensemble in every case considered.
System Gibbs Conventional Boltzmann Modified Boltzmann Canonical
3D ideal gas $E=\frac{3N}{2}k_{B}T$ $E=\left(\frac{3N}{2}-1\right)k_{B}T$ $E=\frac{3N}{2}k_{B}T$ $E=\frac{3N}{2}k_{B}T$
Harmonic oscillator (high $T$) $E=k_{B}T-\frac{h\nu}{2}$ $T=\infty$ $E=k_{B}T$ $E=k_{B}T$
Harmonic oscillator (any $T$) [^1] Note that $E_{0}=\frac{1}{2}h\nu$ and $\theta=\frac{h\nu}{k_{B}}$ for the harmonic oscillator. $\frac{E-E_{0}}{h\nu}=\frac{3-\exp\left(\frac{\theta}{T}\right)}{2\exp\left(\frac{\theta}{T}\right)-2}$ $T=\infty$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T}\right)-1}$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T}\right)-1}$
Particle in 1D box $E=\frac{k_{B}T}{2}$ $E=-\frac{k_{B}T}{2}$ $E=\frac{k_{B}T}{2}$ $E=\frac{k_{B}T}{2}$
Negative temperatures? No Yes Yes Yes
Application of the modified entropy to example systems: $d$-dimensional monoatomic ideal gas. Consider a $d$-dimensional monoatomic ideal gas consisting of $N$ atoms of mass $m$ each, occupying a total volume $V$. This system has unbounded energy levels and thus should not have negative temperature, as correctly argued by Dunkel and Hilbert[5]. The volume of the phase space for this system is [19]
$\Omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}+1\right)}E^{\frac{dN}{2}}.$
Further, the number of states can be written as
$\omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}-1}\epsilon.$
The Gibbs temperature ($T_{G}$) can be calculated using the equation $\frac{1}{T_{G}}=\frac{\partial S_{G}}{\partial E}$ to obtain $E=\frac{dN}{2}k_{B}T_{G}$. However, using the analogous equation for Boltzmann temperature, assuming a fixed energy tolerance $\epsilon$, leads to $E=\left(\frac{dN}{2}-1\right)k_{B}T_{B}$. Note that if $dN=1$, i.e., the system has one degree of freedom, one obtains negative temperatures whereas if $dN=2$, i.e., the system has two degrees of freedom, one gets infinite temperature [5]. This is a serious shortcoming of the Boltzmann entropy, as pointed out by Dunkel and Hilbert [5], as one is led to predict negative/infinite temperature for a system whose energy is unbounded. Considering a fractional energy tolerance, one however obtains $\omega_{r}=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}}\epsilon_{r}$, which combined with Eq. (6) leads to
$E=\frac{dN}{2}k_{B}T_{B_{r}}.$
The above equation not only disallows negative/infinite temperatures for an ideal gas but fixes the disagreement between the energy of a three-dimensional ideal gas as traditionally calculated in the canonical $\left(E=\frac{3N}{2}k_{B}T\right)$ and microcanonical $\left(E=\left(\frac{3N}{2}-1\right)k_{B}T\right)$ ensembles (see Table 1)!
Quantum harmonic oscillator. For a quantum harmonic oscillator, the energy of the system is given as [40] $E_{n}=h\nu\left(n+\frac{1}{2}\right)$, where $n=0,1,2,...$ and $\nu$ represents the frequency of vibration of the oscillator. It follows that $\Omega=1+n=\frac{E}{h\nu}+\frac{1}{2}$. At the high-temperature limit, the states can be treated as a continuum, so that $k_{B}T_{G}=\frac{h\nu}{2}+E$. Although applying the original definition of Boltzmann entropy leads to infinite temperature, the newly proposed definition (Eq. (5)) results in $\omega_{r}=\frac{\epsilon_{r}E}{h\nu}$, leading to $E=k_{B}T_{B_{r}}$, which predicts finite, positive temperatures. Again, only the relative-energy-constrained Boltzmann entropy leads to predictions consistent with the canonical ensemble (Table 1)! One can also derive expressions valid for the temperature under all cases, using discretized definitions of the partition function and temperature (see, e.g., ref. [41]), as shown in the Supplementary Information [42] and in Table 1. The ensuing $E$ vs. $T$ relationships for various entropy choices are plotted in Figure 2. Therein, one sees that only the modified Boltzmann entropy predicts thermodynamically consistent results at all temperatures, including the existence of the zero-point energy $\left(\frac{1}{2}h\nu\right)$ and agreeing with the high-temperature prediction of $E=k_{B}T$. On the contrary, while the Gibbs entropy fails to predict the existence of a zero-point energy (i.e., leads to a Planck oscillator rather than a Schrodinger oscillator [14]!), the Boltzmann entropy predicts infinite temperatures for all energies. Note that, to the best of the author’s knowledge, this is the first derivation of the temperature of a single, isolated quantum harmonic oscillator, i.e., in the microcanonical ensemble; previous expressions relied on the consideration of multiple oscillators, in the limit of a large number of oscillators (see, e.g., [43] and [44]).
Figure 2: Comparison of various microcanonical entropies in terms of their ability to predict the thermodynamic behavior of a quantum harmonic oscillator. The canonical prediction is shown using a blue line and the high-temperature canonical/modified Boltzmann predictions are shown using a dashed black line. Blue, red, and green circles represent predictions of the modified Boltzmann, Gibbs, and conventional Boltzmann entropies in the microcanonical ensemble. In these cases, lines are simply a guide. Only the modified Boltzmann entropy predicts the existence of the correct zero-point energy at 0 K and degenerates to the limit of $k_{B}T$ at high temperatures.
Quantum particle in a 1D box. The energy of a quantum particle in a one-dimensional (1D) box is given as [40] $E_{n}=\frac{an^{2}}{L^{2}}$, where $a=\frac{\hbar^{2}\pi^{2}}{2m}$ and $n=1,2,3,...$, so that $\Omega=L\sqrt{\frac{E}{a}}$. At high temperatures, the states form a continuum, and we find that $E=\frac{k_{B}T_{G}}{2}$ and $P_{G}=T_{G}\frac{\partial S_{G}}{\partial L}=\frac{2E}{L}$. This statistical-mechanical pressure equals the thermodynamic pressure calculated as $P_{T}=-\frac{\partial E}{\partial L}$. Now, although the conventional Boltzmann entropy leads to a negative temperature $\left(E=-\frac{k_{B}T_{B}}{2}\right)$ and negative pressure $\left(P_{B}=-\frac{2E}{L}\right)$, the newly proposed definition of Boltzmann entropy fixes these predictions, as seen in Table 1, to yield $E=\frac{k_{B}T_{B_{r}}}{2}$ and $P_{B_{r}}=\frac{2E}{L}$.
Two-level system and the possibility of negative temperatures. We now demonstrate that the newly proposed definition of Boltzmann entropy still allows for negative temperatures for systems having non-monotonic density of states [21, 22, 12], i.e., a bounded phase space. Considering the well-known case of a two-state system, with energy levels 0 and $E_{1}$ and $N_{0}$ particles occupying the ground state and $N_{1}$ particles occupying the excited state, the energy of the system is $E=N_{1}E_{1}$ and the number of particles is $N=N_{0}+N_{1}$. It follows that the number of states with energy less than or equal to $E$ is
$\Omega=\sum_{N_{0}^{\prime}=1}^{N_{1}}\frac{N!}{N_{0}^{\prime}!N_{1}^{\prime}!}$
where $N_{1}=E/E_{1}$ and $N_{1}^{\prime}=N-N_{0}^{\prime}$. The original Boltzmann partition function is given as
$\omega=\frac{\Delta\Omega}{\Delta E}\epsilon=\frac{N!}{(N_{0}-1)!(N_{1}+1)!}\frac{\epsilon}{E_{1}}.$
where $\Delta\Omega$ is obtained as the difference in the summation expressions of $\Omega$ with $N_{0}^{\prime}$ going up to $(N_{1}+1)$ and $N_{1}$. Considering a large number of particles and applying Stirling’s approximation results in
$S_{B}=k_{B}N\ln N-k_{B}(N_{0}-1)\ln(N_{0}-1)-$
$k_{B}(N_{1}+1)\ln(N_{1}+1)+k_{B}\ln\frac{\epsilon}{E_{1}}.\;\;\;\;\;\;\;$
For fixed $N$, the absolute-energy-constrained Boltzmann temperature can be determined to be
$T_{B}=\frac{\Delta E}{\Delta S_{B}}=\frac{E_{1}}{k_{B}\ln\left(\frac{(N_{0}-1)^{N_{0}-1}(N_{1}+1)^{N_{1}+1}}{(N_{0}-2)^{N_{0}-2}(N_{1}+2)^{N_{1}+2}}\right)}.$
Since we are considering a large number of particles, the temperature reduces to $T_{B}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$. Now, if $E_{1}>0$ and $N_{0}<N_{1}$, i.e., we consider a system exhibiting “population inversion”, we find that $T_{B}<0$. Analogously, the modified Boltzmann partition function is given as
$\omega_{r}=\frac{\Delta\Omega}{\Delta E}\epsilon_{r}E=\frac{N!}{(N_{0}-1)!(N_{1}+1)!}\frac{\epsilon_{r}E}{E_{1}}.$
This leads to, via the Stirling’s approximation, the following expression for the temperature:
$T_{B_{r}}=\frac{\Delta E}{\Delta S_{B_{r}}}=\frac{E_{1}}{k_{B}\ln\left(\frac{(N_{0}-1)^{N_{0}-1}(N_{1}+1)^{N_{1}+2}}{(N_{0}-2)^{N_{0}-2}(N_{1}+2)^{N_{1}+2}N_{1}}\right)},$
which also simplifies, for a large number of particles, to $T_{B_{r}}=\frac{E_{1}}{k_{B}\ln\left(\frac{N_{0}}{N_{1}}\right)}$, also allowing negative temperatures. (This is the same expression that one obtains by direct enumeration as well [12].) On the other hand, Gibbs entropy would not admit negative temperatures at all due to a monotonically increasing total number of states. Thus, Gibbs entropy cannot account for systems with bounded spectra in a thermodynamically consistent manner [45, 46, 12]. Note that this also indicates that Gibbs and Boltzmann entropies can disagree, even in the thermodynamic limit. Now, according to the above discussion, the newly proposed relative-energy-constrained Boltzmann entropy ($S_{B_{r}}$) allows for both negative and positive temperatures, and in the appropriate contexts in which they should be seen, i.e., systems with bounded and unbounded phase space, respectively (see Table 1). Therefore, this work indicates that previous negative temperature measurements for various systems are indeed correct [47, 48, 49, 50, 51, 52, 53], unlike what is stated in ref. [5]. In fact, another argument against negative temperature, viz., the incorrect interpretation of Carnot efficiencies being more than one has also been addressed by Abraham and Penrose [54].
Conclusions. In this work, we introduced a relative energy constraint in the calculation of the microcanonical partition function, which eliminates negative/infinite temperatures for systems with unbounded spectra, such as ideal gases with a few degrees of freedom, a quantum harmonic oscillator, and a quantum particle confined in an infinite well. At the same time, the proposed entropy definition allows for negative temperatures for systems with non-monotonic density of states, i.e., a bounded spectrum, such as a two-state system with population inversion. This makes the proposed entropy definition the one that is the closest to predictions from the canonical ensemble in each case considered (see Table 1). The physical insight underlying the proposed approach is that systems with higher energy will interact more with their surroundings, thus causing a larger perturbation in their energy. Mathematically, while an absolute energy constraint uses up one degree of freedom from the system (by fixing the value of $E$), the use of a relative energy constraint fixes only the ratio $H/E$, thus retaining the original number of degrees of freedom of the system and resulting in the correct scaling of energy with temperature for various systems. Thereby, we proposed a resolution to the ongoing debate in microcanonical statistical mechanics on the use of the Gibbs volume entropy and Boltzmann surface entropy. Overall, this work sets the foundation for using the relative-energy-constrained Boltzmann surface entropy in the microcanonical ensemble of statistical mechanics. Thereby, the work puts on firm theoretical footing the existence of negative temperatures for systems with bounded energy spectra. Finally, this work lays a solid foundation for accurate theoretical descriptions of isolated systems with few degrees of freedom using statistical mechanics.
Acknowledgements.
A.G.R. acknowledges the Infosys Foundation, Bengaluru for an Infosys Young Investigator grant. A.G.R. thanks Profs. B. J. Cherayil, K. G. Ayappa, N. M. Dixit, P. K. Maiti, P. R. Nott, and S. N. Punnathanam for insightful discussions, and Profs. A. Gupta and S. K. Gupta for valuable suggestions regarding the manuscript.